OpenStudy (anonymous):

HELP MEDAL + FAN! 9/36 = ? Write the above ration in its lowest term

OpenStudy (jadedry):

36/9 = ?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

No 9/36 @Jadedry

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Its a Fraction

OpenStudy (jadedry):

@Sabu I know, but 36/9 = 4 Therefore 9/36 = 1/4

OpenStudy (jadedry):

Because 4 also equals 4/1